# New Lower Bounds for Stochastic Non-Convex Optimization through Divergence Composition

### v1

@@ Abstract
We study fundamental limits of first-order stochastic optimization in a range of non-convex settings, including $L$-smooth functions satisfying Quasar-Convexity (QC), Quadratic Growth (QG), and Restricted Secant Inequalities (RSI). While the convergence properties of standard algorithms are well-understood in deterministic regimes, significantly fewer results address the stochastic case, where only unbiased and noisy gradients are available. We establish new lower bounds on the number of noisy gradient queries to minimize these classes of functions, also showing that they are tight (up to a logarithmic factor) in all the relevant quantities characterizing each class. Our approach reformulates the optimization task as a function identification problem, leveraging divergence composition arguments to construct a challenging subclass that leads to sharp lower bounds. Furthermore, we present a specialized algorithm in the one-dimensional setting that achieves faster rates, suggesting that certain dimensional thresholds are intrinsic to the complexity of non-convex stochastic optimization.
Keywords: Non-convex optimization, stochastic first-order optimization, oracle complexity, minimax lower bounds.
@@ 1 Overview and Main Contributions
For the class of convex functions, lower bounds are established using quadratic and piece-wise linear functions that ensure a relative entropy between the distributions of oracle feedbacks at a given point that is uniform (independent of the queried point) (Agarwal et al., 2009; Raginsky and Rakhlin, 2011). Such constructions suffice to obtain sharp lower bounds for convex functions.
In contrast, the function classes we consider here offer greater flexibility in constructing a “difficult” subclass of functions. Specifically, the non-convexity allows us to create functions whose curvature varies across different regions, leading to a relative entropy between the oracle feedback distributions that depend on where the query is made. We leverage this by designing functions that exhibit large curvature locally around the minimizer and smaller curvature away from it. Such a construction guarantees that misidentifying the function incurs a substantial error due to the large local curvature, while the overall relative entropy between feedback distributions remains limited because the algorithm must query near the minimizer (unknown a priori) in order to observe a higher relative entropy compared to other regions. This intuition is captured by the following result, known as divergence composition in the bandits/active learning literature (Lattimore and Szepesvári, 2020). To the best of our knowledge, this technique has not been previously leveraged in the analysis of stochastic first-order optimization, and its application here provides a novel way to derive lower bounds.
Divergence Composition.
@@
We tailor the above analysis to different classes of non-convex functions, yielding the following new results. Throughout, $\sigma^{2}$ denotes the bound on the variance of the stochastic gradients provided by the oracle.
@@
(Informal) Let $\mathcal{X}=B_{d}(\bm{0},D)$. Suppose that $d=\Omega\left(\log(2/\tau)\right)$ and $T$ exceeds a threshold depending on $\sigma,L,$ and $D$. For any optimization algorithm, there exists an $L$-smooth, $\tau$-QC function $f$ such that
@@
(Informal) Let $\mathcal{X}=\mathbb{R}^{d}$. Suppose that $d=\Omega\left(\log(2/\tau)\right)$. For any optimization algorithm, there exists an $L$-smooth, $\tau$-QC, and $\mu$-QG function $f$ such that
@@
(Informal) Let $\mathcal{X}=\mathbb{R}^{d}$ and $\kappa=L/\mu$. Suppose that $d=\Omega\left(\log(2\kappa)\right)$. For any optimization algorithm, there exists an $L$-smooth, $\mu$-RSI function $f$ such that
@@
Finally, we present an argument indicating that the necessity for $d$ to exceed a certain problem-dependent threshold is an intrinsic characteristic of the optimization problem itself rather than merely a byproduct of our analysis. Specifically, in Section 4, we propose a novel algorithm tailored for the one-dimensional setting and demonstrate the following better upper bound.
@@ 2 Related Work and Classes Comparisons
Various classes of non-convex functions have received attention in the optimization literature. Notable examples include the Polyak-Łojasiewicz (PL) condition (Polyak, 1963), Error Bounds (EB) (Luo and Tseng, 1993), Quadratic Growth (QG) (Anitescu, 2000), Essential Strong Convexity (ESC) (Liu et al., 2014), Restricted Secant Inequality (RSI) (Polyak, 1963), Restricted Strong Convexity (RSC) (Zhang and Yin, 2013), and Quasi-Strong Convexity (Necoara et al., 2019). All these classes can be seen as gradual weakening of the strong convexity (SC) assumption, till including non-convex functions (Karimi et al., 2016).
@@
Most of the previous work on both upper and lower bounds has focused on the deterministic setting, i.e., when the gradients are exact, or on in the finite-sum setting.
@@ 3 Lower Bounds
Let $\mathcal{F}$ be a class of objective functions. An algorithm $\mathcal{A}$ solving the problem (1) interacts iteratively with a stochastic first-order oracle. At each iteration $t=1,\dots,T$, the algorithm selects a query point $\bm{x}_{t}\in\mathcal{X}$ and receives an oracle response $\phi(\bm{x}_{t},f)$, which represents a noisy gradient of $f$ at $\bm{x}_{t}$. The noise is assumed to be zero-mean with bounded variance or sub-Gaussian, depending on the oracle considered. Based on the sequence of past responses, $\{\phi(\bm{x}_{1},f),\dots,\phi(\bm{x}_{t},f)\}$, the algorithm determines the next query point $\bm{x}_{t+1}$. Our analysis focuses on the information-theoretic complexity of this procedure, aiming to determine the number of oracle queries required to achieve a given accuracy for any function in $\mathcal{F}$.
@@
Within this class, we say that $\phi$ is a sub-Gaussian oracle if, additionally, the random vector $\phi(\bm{x},f)$ is $\sigma$-sub-Gaussian.
We denote $\mathbb{O}_{\sigma}$ the class of first-order stochastic oracles and $\mathbb{O}^{\mathrm{sg}}_{\sigma}$ the class of first-order $\sigma$-sub-Gaussian. Note that we have $\mathbb{O}^{\mathrm{sg}}_{\sigma}\subset\mathbb{O}_{\sigma}$. Therefore a lower bound that holds over the class $\mathbb{O}^{\mathrm{sg}}_{\sigma}$ holds over the class $\mathbb{O}_{\sigma}$. We state below the lower bounds for the classes considered.
@@
$\sup_{\phi\in\mathbb{O}_{\sigma}^{\mathrm{sg}}}\ \epsilon^{*}_{T}(\mathrm{QC}\cap\mathrm{QG},\mathcal{X};\phi)\geq c\cdot\frac{\sigma^{2}}{\mu\tau^{2}\log(2/\tau)T}~{}.$
@@
$\sup_{\phi\in\mathbb{O}_{\sigma}^{\mathrm{sg}}}\ \epsilon^{*}_{T}(\mathrm{RSI},\mathcal{X};\phi)\geq c\cdot\frac{L\sigma^{2}}{\mu^{2}\log(2\kappa)T}~{}.$
@@
$\sup_{\phi\in\mathbb{O}_{\sigma}^{\mathrm{sg}}}\ \epsilon^{*}_{T}(\mathrm{QC},\mathcal{X};\phi)\geq c\cdot\frac{D\sigma}{\tau\sqrt{\ln(2/\tau)T}}~{}.$
@@
We introduce the following additional notation. For any integer $n\geq 1$, let $\mathcal{F}_{n}$ denote the class of real-valued functions on $\mathbb{R}^{n}$ that are $L$-smooth, $\tau-\mathrm{QC}$, and satisfy the $\mu-\mathrm{QG}$. Let $A:=\{1,\dots,d_{0}\}$ and $\bar{A}:=[d]\setminus A$. For any $\bm{x}\in\mathbb{R}^{d}$, let $\bm{x}_{A}\in\mathbb{R}^{d_{0}}$ be the vector of its first $d_{0}$ coordinates, and $\bm{x}_{\bar{A}}\in\mathbb{R}^{d-d_{0}}$ be the vector of its remaining coordinates. For two vectors $\mathbf{a}\in\mathbb{R}^{d_{0}}$ and $\mathbf{b}\in\mathbb{R}^{d-d_{0}}$, we write $(\mathbf{a},\mathbf{b})$ to denote the vector in $\mathbb{R}^{d}$ whose first $d_{0}$ components are $\mathbf{a}$ and last $d-d_{0}$ components are $\mathbf{b}$.
@@
Oracle Construction. We now specify the stochastic oracle $\phi:\mathbb{R}^{d}\times\mathcal{F}_{d}\to\mathbb{R}^{d}$ used in the proof. Let $\bm{\xi}$ be a sample from a $d_{0}$-dimensional normal distribution with zero mean and covariance matrix $\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}$, i.e., $\bm{\xi}\sim\mathcal{N}_{d_{0}}\bigl{(}\bm{0}_{A},\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\bigr{)}$. Given an input $\bm{x}\in\mathbb{R}^{d}$ and a function $H\in\mathcal{F}_{d}$, we define $\phi(\bm{x},H)=\nabla H(\bm{x})+(\bm{\xi},\mathbf{0}_{\bar{A}})$. Thus, $\phi$ is a $\sigma$-sub-Gaussian stochastic oracle as specified in Definition 3.1. Moreover, for each $i\in[m]$, we have $\phi(\bm{x},F_{i})=\bigl{(}\nabla f_{i}(\bm{x}_{A})+\bm{\xi},\bm{0}_{\bar{A}}\bigr{)}$.
Divergence Composition. We reduce the optimization problem to one of function identification. To that end, consider a “reference function” $G:\mathbb{R}^{d}\to\mathbb{R}$ that also depends only on its first $d_{0}$ coordinates. In other words, there exists a function $g:\mathbb{R}^{d_{0}}\to\mathbb{R}$ such that for every $\bm{x}\in\mathbb{R}^{d}$, $G(\bm{x})=g(\bm{x}_{A})$. We choose $g$ so that its gradient satisfies, for any $\bm{x}\in\mathbb{R}^{d_{0}}$,
@@
To derive an upper bound on $\mathbb{P}(\mathcal{E}_{i})$ (or equivalently, a lower bound on the probability of $\neg\mathcal{E}_{i}$), we apply Pinsker’s inequality. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$
@@ 5 Conclusion
In this work, we establish new lower bounds for first-order stochastic optimization for several non-convex function classes. Specifically, we derive lower bounds on the number of noisy gradient queries necessary to minimize $L$-smooth functions satisfying quasar-convexity plus quadratic growth or restricted secant inequalities. Our approach leverages a divergence composition technique, allowing us to construct hard instances that lead to nearly tight lower bounds on optimization complexity.
@@ A.1 Definitions of Classes
    $f(\bm{y})\geq f(\bm{x})+\langle\nabla f(\bm{x}),\bm{y}-\bm{x}\rangle+\frac{\mu}{2}\left\|{y-x}\right\|^{2}~{}.$
@@
    $f(\bm{x}_{p})\geq f(\bm{x})+\langle\nabla f(\bm{x}),\bm{x}_{p}-\bm{x}\rangle+\frac{\mu}{2}\left\|{\bm{x}_{p}-x}\right\|^{2},$
    where $\bm{x}_{p}$ is $\text{Proj}_{S}(x)$.
@@
    where $\bm{x}_{p}$ is $\text{Proj}_{S}(x)$.
@@
    $\frac{1}{2}\left\|{\nabla f(x)}\right\|^{2}\geq\mu(f(\bm{x})-f(x_{p}))~{}.$
@@
    $f(\bm{x})-f^{*}\leq\frac{1}{\tau}\langle\nabla f(\bm{x}),\bm{x}-\bm{x}_{p}\rangle-\frac{\mu}{2}\left\|{x_{p}-x}\right\|^{2},$
    where $\bm{x}_{p}$ is $\text{Proj}_{S}(x)$.
@@
    $\left\|{\nabla f(\bm{x})}\right\|\left\|{\bm{x}-\bm{x}_{p}}\right\|\geq\langle\nabla f(x),x-x_{p}\rangle\geq\mu\left\|{\bm{x}-\bm{x}_{p}}\right\|^{2}~{}.$
@@
Oracle Construction. We now specify the stochastic oracle $\phi:\mathbb{R}^{d}\times\mathcal{F}_{d}\to\mathbb{R}^{d}$ used in the proof. Let $\bm{\xi}$ be a sample from a $d_{0}$-dimensional normal distribution with zero mean and covariance matrix $\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}$, i.e., $\bm{\xi}\sim\mathcal{N}_{d_{0}}\bigl{(}\bm{0}_{A},\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\bigr{)}$. Given an input $\bm{x}\in\mathbb{R}^{d}$ and a function $H\in\mathcal{F}_{d}$, we define $\phi(\bm{x},H)=\nabla H(\bm{x})+(\bm{\xi},\mathbf{0}_{\bar{A}})$. Thus, $\phi$ is a $\sigma$-sub-Gaussian stochastic oracle as specified in Definition 3.1. Moreover, for each $i\in[m]$, we have $\phi(\bm{x},F_{i})=\bigl{(}\nabla f_{i}(\bm{x}_{A})+\bm{\xi},\bm{0}_{\bar{A}}\bigr{)}$.
@@
To derive an upper bound on $\mathbb{P}(\mathcal{E}_{i})$ (or equivalently, a lower bound on the probability of $\neg\mathcal{E}_{i}$), we apply Pinsker’s inequality. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$
@@
We use Pinsker’s inequality to derive an upper bound on $\mathbb{P}(\mathcal{E}_{i})$. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$
@@
We use Pinsker’s inequality to derive an upper bound on $\mathbb{P}(\mathcal{E}_{i})$. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$

### v2

# New Lower Bounds for Stochastic Non-Convex Optimization through Divergence Decomposition

@@ Abstract
We study fundamental limits of first-order stochastic optimization in a range of non-convex settings, including $L$-smooth functions satisfying Quasar-Convexity (QC), Quadratic Growth (QG), and Restricted Secant Inequalities (RSI). While the convergence properties of standard algorithms are well-understood in deterministic regimes, significantly fewer results address the stochastic case, where only unbiased and noisy gradients are available. We establish new lower bounds on the number of noisy gradient queries to minimize these classes of functions, also showing that they are tight (up to a logarithmic factor) in all the relevant quantities characterizing each class. Our approach reformulates the optimization task as a function identification problem, leveraging divergence decomposition arguments to construct a challenging subclass that leads to sharp lower bounds. Furthermore, we present a specialized algorithm in the one-dimensional setting that achieves faster rates, suggesting that certain dimensional thresholds are intrinsic to the complexity of non-convex stochastic optimization.
Keywords: Non-convex optimization, stochastic first-order optimization, oracle complexity, minimax lower bounds
@@ 1 Overview and Main Contributions
For the class of convex functions, lower bounds are established using quadratic and piece-wise linear functions that ensure a relative entropy between the distributions of oracle feedbacks at a given point that is uniform (independent of the queried point) (Nemirovskij and Yudin, 1983; Agarwal et al., 2009; Raginsky and Rakhlin, 2011). Such constructions suffice to obtain sharp lower bounds for convex functions.
In contrast, the function classes we consider here offer greater flexibility in constructing a “difficult” subclass of functions. Specifically, the non-convexity allows us to create functions whose curvature varies across different regions, leading to a relative entropy between the oracle feedback distributions that depend on where the query is made. We leverage this by designing functions that exhibit large curvature locally around the minimizer and smaller curvature away from it. Such a construction guarantees that misidentifying the function incurs a substantial error due to the large local curvature, while the overall relative entropy between feedback distributions remains limited because the algorithm must query near the minimizer (unknown a priori) in order to observe a higher relative entropy compared to other regions. This intuition is captured by the following result, known as divergence decomposition in the bandits/active learning literature (Lattimore and Szepesvári, 2020). To the best of our knowledge, this technique has not been previously leveraged in the analysis of stochastic first-order optimization, and its application here provides a novel way to derive lower bounds.
Divergence Decomposition.
@@
We tailor the above analysis to different classes of non-convex functions, yielding the following new results. Throughout, $\sigma^{2}$ denotes the bound on the variance of the stochastic gradients provided by the oracle. The bounds apply to any randomized algorithm whose query at round $t$ can depend on all past observations and the algorithm’s independent internal randomness. The reported solution $\hat{\bm{x}}$ may be any measurable function of the complete history and need not coincide with the final iterate.
@@
(Informal) Let $\mathcal{X}=B_{d}(\bm{0},D)$. Suppose that $d=\Omega\left(\log(2/\tau)\right)$ and $T$ exceeds a threshold depending on $\sigma,L,$ and $D$. For any optimization algorithm having access to unbiased stochastic gradient with variance bounded by $\sigma^{2}$, there exists an $L$-smooth, $\tau$-QC function $f$ such that
@@
(Informal) Let $\mathcal{X}=\mathbb{R}^{d}$. Suppose that $d=\Omega\left(\log(2/\tau)\right)$. For any optimization algorithm having access to unbiased stochastic gradient with variance bounded by $\sigma^{2}$, there exists an $L$-smooth, $\tau$-QC, and $\mu$-QG function $f$ such that
@@
(Informal) Let $\mathcal{X}=\mathbb{R}^{d}$ and $\kappa=L/\mu$. Suppose that $d=\Omega\left(\log(2\kappa)\right)$. For any optimization algorithm having access to unbiased stochastic gradient with variance bounded by $\sigma^{2}$, there exists an $L$-smooth, $\mu$-RSI function $f$ such that
@@
Finally, we present an argument indicating that the necessity for $d$ to exceed a certain problem-dependent threshold is an intrinsic characteristic of the optimization problem itself rather than merely a byproduct of our analysis. Specifically, in Section 4, we propose a novel algorithm tailored for the one-dimensional setting and demonstrate the following better upper bound. While previous lower-bound constructions for locating stationary points in non-convex smooth functions also required high dimensionality (Arjevani et al., 2023), their arguments demand a dimension that grows polynomially with the problem parameters and the time horizon $T$. By contrast, our bound is satisfied once the dimension merely exceeds a logarithmic function of the problem parameters, with no dependence on $T$.
@@ 2 Related Work and Classes Comparisons
Various classes of non-convex functions have received attention in the optimization literature. Notable examples include the Polyak-Łojasiewicz (PL) condition (Polyak, 1963), Error Bounds (EB) (Luo and Tseng, 1993), Quadratic Growth (QG) (Anitescu, 2000), Essential Strong Convexity (ESC) (Liu et al., 2014), Restricted Secant Inequality (RSI) (Polyak, 1963), Restricted Strong Convexity (RSC) (Zhang and Yin, 2013), and Quasi-Strong Convexity (Necoara et al., 2019). All these classes can be seen as gradual weakening of the strong convexity (SC) assumption, until including non-convex functions (Karimi et al., 2016).
@@
Most of the previous work on both upper and lower bounds has focused on the deterministic setting, i.e., when the gradients are exact, or in the finite-sum setting.
@@
A separate line of research, represented by Arjevani et al. (2023), derives lower bounds for smooth stochastic non-convex problems with the goal of reaching an $\epsilon$-stationary point. In contrast, our work focuses on minimizing the optimization error itself by seeking a global minimizers within the function classes we study.
@@ 3 Lower Bounds
Let $\mathcal{F}$ be a class of objective functions. An algorithm $\mathcal{A}$ solving the problem (1) interacts iteratively with a stochastic first-order oracle. At each iteration $t=1,\dots,T$, the algorithm selects a query point $\bm{x}_{t}\in\mathcal{X}$ and receives an oracle response $\phi(\bm{x}_{t},f)$, which represents a noisy gradient of $f$ at $\bm{x}_{t}$. The noise is assumed to be zero-mean with bounded variance. Based on the sequence of past responses, $\{\phi(\bm{x}_{1},f),\dots,\phi(\bm{x}_{t},f)\}$, the algorithm determines the next query point $\bm{x}_{t+1}$. After $T$ rounds the algorithm outputs $\hat{\bm{x}}_{T}\in\mathcal{X}$. This output is not restricted to the final iterate; it may be any function of the observed history $\{\phi(\bm{x}_{1},f),\dots,\phi(\bm{x}_{T},f)\}$ to the set $\mathcal{X}$. Our analysis focuses on the information-theoretic complexity of this procedure, aiming to determine the number of oracle queries required to achieve a given accuracy for any function in $\mathcal{F}$.
@@
Let $\mathbb{O}_{\sigma}$ denote the family of first-order stochastic oracles introduced in Definition 3.1. In all forthcoming lower-bound results, the learner observes a multivariate Gaussian vector with independent components whose total variance does not exceed $\sigma^{2}$. The corresponding lower bounds for each problem class are stated below.
@@
$\sup_{\phi\in\mathbb{O}_{\sigma}}\ \epsilon^{*}_{T}(\mathrm{QC}\cap\mathrm{QG},\mathcal{X};\phi)\geq c\cdot\frac{\sigma^{2}}{\mu\tau^{2}\log(2/\tau)T}~{}.$
@@
$\sup_{\phi\in\mathbb{O}_{\sigma}}\ \epsilon^{*}_{T}(\mathrm{RSI},\mathcal{X};\phi)\geq c\cdot\frac{L\sigma^{2}}{\mu^{2}\log(2\kappa)T}~{}.$
@@
$\sup_{\phi\in\mathbb{O}_{\sigma}}\ \epsilon^{*}_{T}(\mathrm{QC},\mathcal{X};\phi)\geq c\cdot\frac{D\sigma}{\tau\sqrt{\ln(2/\tau)T}}~{}.$
@@
We introduce the following additional notation. For any integer $n\geq 1$, let $\mathcal{F}_{d}$ denote the class of real-valued functions on $\mathbb{R}^{d}$ that are $L$-smooth, $\tau-\mathrm{QC}$, and satisfy the $\mu-\mathrm{QG}$. Let $A:=\{1,\dots,d_{0}\}$ and $\bar{A}:=[d]\setminus A$. For any $\bm{x}\in\mathbb{R}^{d}$, let $\bm{x}_{A}\in\mathbb{R}^{d_{0}}$ be the vector of its first $d_{0}$ coordinates, and $\bm{x}_{\bar{A}}\in\mathbb{R}^{d-d_{0}}$ be the vector of its remaining coordinates. For two vectors $\mathbf{a}\in\mathbb{R}^{d_{0}}$ and $\mathbf{b}\in\mathbb{R}^{d-d_{0}}$, we write $(\mathbf{a},\mathbf{b})$ to denote the vector in $\mathbb{R}^{d}$ whose first $d_{0}$ components are $\mathbf{a}$ and last $d-d_{0}$ components are $\mathbf{b}$.
@@
Oracle Construction. We now specify the stochastic oracle $\phi:\mathbb{R}^{d}\times\mathcal{F}_{d}\to\mathbb{R}^{d}$ used in the proof. Let $\bm{\xi}$ be a sample from a $d_{0}$-dimensional normal distribution with zero mean and covariance matrix $\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}$, i.e., $\bm{\xi}\sim\mathcal{N}_{d_{0}}\bigl{(}\bm{0}_{A},\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\bigr{)}$. Given an input $\bm{x}\in\mathbb{R}^{d}$ and a function $H\in\mathcal{F}_{d}$, we define $\phi(\bm{x},H)=\nabla H(\bm{x})+(\bm{\xi},\mathbf{0}_{\bar{A}})$. Thus, $\phi$ is a stochastic oracle belonging to $\mathbb{O}_{\sigma}$ as specified in Definition 3.1. Moreover, for each $i\in[m]$, we have $\phi(\bm{x},F_{i})=\bigl{(}\nabla f_{i}(\bm{x}_{A})+\bm{\xi},\bm{0}_{\bar{A}}\bigr{)}$.
Divergence Decomposition. We reduce the optimization problem to one of function identification. To that end, consider a “reference function” $G:\mathbb{R}^{d}\to\mathbb{R}$ that also depends only on its first $d_{0}$ coordinates. In other words, there exists a function $g:\mathbb{R}^{d_{0}}\to\mathbb{R}$ such that for every $\bm{x}\in\mathbb{R}^{d}$, $G(\bm{x})=g(\bm{x}_{A})$. We choose $g$ so that its gradient satisfies, for any $\bm{x}\in\mathbb{R}^{d_{0}}$,
@@
To derive an upper bound on $\mathbb{P}_{i}(\mathcal{E}_{i})$ (or equivalently, a lower bound on the probability of $\neg\mathcal{E}_{i}$), we apply Pinsker’s inequality. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}_{i}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$
@@ 5 Conclusion
In this work, we establish new lower bounds for first-order stochastic optimization for several non-convex function classes. Specifically, we derive lower bounds on the number of noisy gradient queries necessary to minimize $L$-smooth functions satisfying quasar-convexity plus quadratic growth or restricted secant inequalities. Our approach leverages a divergence decomposition technique, allowing us to construct hard instances that lead to nearly tight lower bounds on optimization complexity.
@@ A.1 Definitions of Classes
    $f(\bm{y})\geq f(\bm{x})+\langle\nabla f(\bm{x}),\bm{y}-\bm{x}\rangle+\frac{\mu}{2}\left\|{\bm{y}-\bm{x}}\right\|^{2}~{}.$
@@
    $f(\bm{x}_{p})\geq f(\bm{x})+\langle\nabla f(\bm{x}),\bm{x}_{p}-\bm{x}\rangle+\frac{\mu}{2}\left\|{\bm{x}_{p}-\bm{x}}\right\|^{2},$
    where $\bm{x}_{p}$ is $\text{Proj}_{S}(\bm{x})$.
@@
    where $\bm{x}_{p}$ is $\text{Proj}_{S}(\bm{x})$.
@@
    $\frac{1}{2}\left\|{\nabla f(\bm{x})}\right\|^{2}\geq\mu(f(\bm{x})-f(\bm{x}_{p}))~{}.$
@@
    $f(\bm{x})-f^{*}\leq\frac{1}{\tau}\langle\nabla f(\bm{x}),\bm{x}-\bm{x}_{p}\rangle-\frac{\mu}{2}\left\|{\bm{x}_{p}-\bm{x}}\right\|^{2},$
    where $\bm{x}_{p}$ is $\text{Proj}_{S}(\bm{x})$.
@@
    $\left\|{\nabla f(\bm{x})}\right\|\left\|{\bm{x}-\bm{x}_{p}}\right\|\geq\langle\nabla f(\bm{x}),\bm{x}-\bm{x}_{p}\rangle\geq\mu\left\|{\bm{x}-\bm{x}_{p}}\right\|^{2}~{}.$
@@
Oracle Construction. We now specify the stochastic oracle $\phi:\mathbb{R}^{d}\times\mathcal{F}_{d}\to\mathbb{R}^{d}$ used in the proof. Let $\bm{\xi}$ be a sample from a $d_{0}$-dimensional normal distribution with zero mean and covariance matrix $\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}$, i.e., $\bm{\xi}\sim\mathcal{N}_{d_{0}}\bigl{(}\bm{0}_{A},\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\bigr{)}$. Given an input $\bm{x}\in\mathbb{R}^{d}$ and a function $H\in\mathcal{F}_{d}$, we define $\phi(\bm{x},H)=\nabla H(\bm{x})+(\bm{\xi},\mathbf{0}_{\bar{A}})$. Thus, $\phi$ is a stochastic oracle belonging to $\mathbb{O}_{\sigma}$ as specified in Definition 3.1. Moreover, for each $i\in[m]$, we have $\phi(\bm{x},F_{i})=\bigl{(}\nabla f_{i}(\bm{x}_{A})+\bm{\xi},\bm{0}_{\bar{A}}\bigr{)}$.
@@
To derive an upper bound on $\mathbb{P}_{i}(\mathcal{E}_{i})$ (or equivalently, a lower bound on the probability of $\neg\mathcal{E}_{i}$), we apply Pinsker’s inequality. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}_{i}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$
@@
We use Pinsker’s inequality to derive an upper bound on $\mathbb{P}_{i}(\mathcal{E}_{i})$. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}_{i}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$
@@
We use Pinsker’s inequality to derive an upper bound on $\mathbb{P}_{i}(\mathcal{E}_{i})$. This gives
$\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}_{i}(\mathcal{E}_{i})\leq\frac{1}{m}\sum_{i=1}^{m}\mathbb{Q}(\mathcal{E}_{i})+\sqrt{\frac{1}{2m}\sum_{i=1}^{m}\operatorname{KL}\left(\mathbb{Q},\mathbb{P}_{i}\right)}~{}.$